Given that M(p,7) is the midpoint of the line segment joining the points S(-3,1) and R(11,q). Find the values of p and q.

Grade:9

1 Answers

Yuvraj
10 years ago
Use mid-point formula: M(x,y)=(x1+x2/2,y1+y2/2) So p=-3+11/2=4 7=1+q/2 so q=13